# Community-level Contagion among Diverse Financial Assets

### v1

@@ 5.2 Intra-community vs Inter-community Contagion
Except for the Covid-19 Outbreak sub-period with 13 communities forming from 221 assets, other sub-periods have at least 15 communities, leading to more than 100 inter-community contagions in each sub-period. Given this huge amount of information, we have to come up with a way to interpret them all. With this in mind, for each sub-period, we construct a distribution of inter-community contagions, thereby the contagion effects between pairs of communities can be understood transparently in each sub-period while being beneficial for further experiments conducted in this study. Additionally, we also construct a distribution for intra-community contagions in each sub-period to facilitate the comparison in intra-community contagion effects across sub-periods, which will be presented subsequently.
We first discuss the changes in contagion effects across the examined sub-periods. Specifically, Figures 5 and 6 display the distributions of intra-community and inter-community contagion metrics (e.g. density (a) and magnitude (b)) across seven sub-periods, respectively. Each sub-period is drawn by a different color. Moreover, we purposely highlight the distributions of some sub-periods that behave relatively differently from the rest using solid lines to help readers better locate them in the figures, while other sub-periods are marked by dash lines. We find that both intra-community and inter-community contagion densities are dramatically amplified during the Covid-19 Outbreak period (solid orange line), compared to other time intervals, as shown in Figures 5(a) and 6(a). This corresponds to one of the most turbulent phases in financial markets, characterized by sharp price declines, heightened volatility, and widespread investor pessimism [79]. This result suggests that the amplified turbulence in financial markets and the extreme reactions of investors (e.g. purchasing and selling herding behavior) caused by the pandemic trigger contagion not only within communities of highly similar assets but also across different asset communities. Furthermore, the Bull Time 2 sub-period (solid red line) also shows a moderate increase in contagion density, although this phenomenon is much weaker than the Covid-19 Outbreak time. This was a period when most assets experienced a price surge, indicating that contagion is not only a feature of negative shocks but can also be triggered under favorable market conditions. In fact, the Covid-19 pandemic had a much stronger impact on financial markets than the bullish time, with widespread systemic disruptions and heightened investor attention [12], explaining for a more pronounced contagion effect observed during that time.
@@
Figure 5: Distributions of a) intra-community contagion densities and b) intra-community contagion magnitudes during each of seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia 1 and Ukraine-Russia 2. We use solid lines to highlight the distributions of specific sub-periods, otherwise, they are drawn by dash lines
@@
Figure 6: Distributions of a) inter-community contagion densities and b) inter-community contagion magnitudes during each of seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia 1 and Ukraine-Russia 2. We use solid lines to highlight the distributions of specific sub-periods, otherwise, they are drawn by dash lines
@@
Remarkably, the distributions of contagion magnitudes across all sub-periods remain relatively similar, indicating a consistent level of magnitude over time, regardless of market conditions. This pattern holds true for both intra-community and inter-community contagions, suggesting that while the number of contagion links tend to increase during turbulent times, the average strength of these links remains stable. This phenomenon seems to be attributed to the portfolio diversification and risk-hedging strategy from investors in which they adjust their holdings across multiple assets rather than only focus on a particular one, especially during a financial crisis [105]. These widespread but diluted approaches can raise the scale of contagion without amplifying the strength of individual spillovers.
@@
We now compare the contagion effects between intra-community and inter-community during each sub-period. Utilizing the distribution of inter-community contagions described above (shown in Figure 6), our idea for this experiment is to compare each intra-community contagion with the distribution to see if the intra-community one behaves similarly to the inter-community contagions or not. In this regard, an intra-community contagion is considered to be significantly stronger than inter-community contagions within a sub-period if its quantile value falls within the right-hand side rejection area in the inter-community contagion distribution, with the significance level being 1% (i.e. the quantile value is greater than or equal to the 99th). Otherwise, it behaves similarly to inter-community contagions. For this, Tables 4-5 show the contagion density and magnitude, respectively, within each community along with their quantiles (inside the brackets) in the corresponding inter-community contagion distribution for each of the seven sub-periods. Generally, both tables reveal that the intra-community contagion effects do not look much different from inter-community contagion effects since most intra-community contagion signals are similar to inter-community contagion signals. Specifically, the density (Table 4) and magnitude (Table 5) values within each community mostly belong to the 99% confidence interval of their corresponding inter-community contagion distribution and are off the rejection areas, indicating that intra-commmunity contagion tends to follow the same behavior as inter-community contagion, across different time intervals and market conditions. By contrast, only a few communities in certain sub-periods reveal an abnormal behavior where their intra-community contagion densities or magnitudes are significantly stronger than the others and are off the inter-community contagion distributions. This phenomenon is more frequently for contagion density. For example, community #9 shows a contagion density exceeding 26% in Bull Time 1 and 32% in Ukraine-Russia Conflict 1, placing it at the 100th percentile of the corresponding inter-community contagion density distributions for those periods. Similarly, communities #7 (100th percentile) and #11 (99th percentile) in Bull Time 2 also exhibit unusually high contagion densities. As for contagion magnitude, only communities #1 and #14 in the Ukraine-Russia Conflict 2 sub-period surpasses inter-community levels. A common characteristic from these communities is that the assets within each community mostly come from the same business sector, either Information Technology, Financials or cryptocurrency. For this result, a possible reason is that these assets are closely relevant to each other, sharing common fundamental drivers such as macroeconomic variables, regulatory changes and investors’ behaviors. This similarity in risk exposures and underlying characteristics naturally leads to higher correlation in price movements, which tends to trigger contagion within the group, especially during market uncertainties [3]. However, these communities only account for a very small part of the network.
@@
Another noteworthy finding is that although cryptocurrencies are highly correlated with each other and exhibit distinct behavior from other traditional assets, as shown in Subsection 5.1, the contagion density and magnitude within communities of cryptocurrencies generally do not differ from other asset groups. The only notable exception is during the Bull Time 2 sub-period, when the contagion density among cryptocurrencies stood the highest level across the entire network, exceeding both intra- and inter-community contagion densities. As discussed earlier, cryptocurrencies share several common fundamental drivers, especially the strong sensitivity of investors to positive market conditions [12], combined with the positive momentum of financial markets during this bullish time, which explains this high contagion density.
In overall, the above results suggest that financial contagion tends to manifest similarly both within and across communities since there is no significant difference between intra-community and inter-community contagions. This implies that contagion is unlikely to be an important factor in the formation of communities. Supporting this point is the case of cryptocurrencies: although they exhibit strong internal correlations and form distinct communities, their contagion levels are generally comparable to those of other asset communities and cross-communities throughout most periods.
@@ 5.3 Directional Inter-community Contagion
In this subsection, we take a closer look at inter-community contagion effects by analyzing how shocks propagate directionally between different communities. Specifically, we investigate how one community transmits shocks to another and how it, in turn, receives shocks. Our goal is to identify whether there exist communities acting as dominant transmitters (e.g. showing strong outgoing contagion to a high number of communities) or receivers (showing strong incoming contagion from a high number of communities). Also, as discussed in the previous subsection, contagion magnitude shows limited variation over time and offers few meaningful insights. Therefore, for the sake of clarity and simplicity, we focus solely on contagion density in this part of the analysis. To this end, we use the directional contagion density metric described in Equation 14 for this experiment.
@@
Figure 7: Directional inter-community contagion density heatmap and its graph representation of top 25% in contagion density.
Indeed, our experimental results show that there are always communities acting as strong transmitters in each sub-period, spreading shocks to many other communities with significantly higher contagion density levels relative to the overall network. On the other hand, this pattern for the case of receiving communities is not observed across all sub-periods, suggesting that there is no community intensively receiving shocks from many others. Notably, while mixed communities (e.g. comprising assets from various sectors) appear in certain cases, the communities that act as strong transmitters in the network are often composed of assets from the Information Technology sector. Besides, Healthcare and Financials sectors also emerge as strong transmitters in specific periods, including the Bull Time 3 for the former and the Ukraine-Russia Conflict 1 for the latter. To support these findings, we present two examples: the graph visualizations of top 25% inter-community contagions in terms of contagion density (Figures 8(a) and 8(b)) and their quantitative details (Tables 6 and 6) for the Bull Time 3 and Ukraine-Russia Conflict 1 sub-periods, respectively. Specifically, Table 6 shows that communities #1 and #2 during the Bull Time 3 sub-period act as dominant transmitters, transmitting high-density shocks (i.e. belong to the top 25% of inter-community contagions in terms of the contagion density) to 14 and 12 out of 16 communities, respectively. These numbers are significantly higher than those for other communities. Moreover, the average contagion densities from communities #1 and #2 to other communities also stand at the highest levels, with 23.75% and 20.10%, respectively . Consequently, these two communities are highlighted in the darkest blue in Figure 8(a). Notably, community #1 is composed predominantly of Information Technology assets (75%) while Healthcare assets contribute approximately 86% to community #2. Likewise, during the Ukraine-Russia Conflict 1 sub-period, communities #6, #9 and #10 emerge as dominant transmitters. For this, Financials assets primarily make up community #9, whereas Information Technology assets account for the majority in the other two communities, further details are provided in Table 6.
@@
Figure 8: Graph representation of the top 25% inter-community contagions in terms of the contagion density during the Bull Time 3 (a) and Ukraine-Russia Conflict 1 (b) sub-periods.
Table 6: Quantitative details of top 25% inter-community contagion links in terms of density during the (a) Bull Time 3 and (b) Ukraine-Russia Conflict 1 sub-periods. Com represents the community, #In is the number of incoming contagion links to a community, #Out is the number of outgoing contagion links from a community, Avg_In is the average contagion density of incoming contagion links to a community, and Avg_Out is the average contagion density of outgoing contagion links from a community. Communities acting as dominant transmitters are highlighted in bold.
@@
The strong influence of these sectors can be attributed to their intrinsic characteristics and also the prevailing market conditions during the examined periods. In particular, Information Technology assets consistently attract significant investor attention due to their strong growth potential and dominant market capitalization in the financial markets, making them highly influential in shaping both investor behavior and overall market performance. As a result, price movements in this sector often have a broad impact across various business sectors and types of investment vehicle. In fact, the sector’s influence has grown even more remarkably since 2020 thanks to the accelerated global technology adoption followed by the boom in generative AI [106, 107].
Regarding the Healthcare and Financials sectors, their strong influence appears to be more context-dependent. During the pandemic, the Healthcare sector gained prominence as companies were at the forefront of efforts to combat Covid-19, including the production of protective equipment and vaccines [108, 109]. This rapid development elevated their importance and impact on both financial markets and the broader economy. Likewise, the Financials sector plays a vital role in reflecting and shaping economic conditions [110], thereby carrying a strong influence in financial markets, especially during periods of uncertainty like the Ukraine-Russia conflict.

### v2

@@ 5.2 Intra-community vs Inter-community Contagion
Except for the Covid-19 Outbreak sub-period with 13 communities forming from 221 assets, other sub-periods have at least 15 communities, leading to more than 100 inter-community contagions in each sub-period. Given this huge amount of information, we have to come up with a way to interpret it all. With this in mind, for each sub-period, we construct a distribution of inter-community contagions, thereby the contagion effects between pairs of communities can be understood transparently in each sub-period while being beneficial for further experiments conducted in this study. Additionally, we also construct a distribution for intra-community contagions in each sub-period to facilitate the comparison of intra-community contagion effects across sub-periods, which will be presented subsequently.
We first discuss the changes in contagion effects across the examined sub-periods. Specifically, Figures 5 and 6 display the distributions of intra-community and inter-community contagion metrics (e.g. density (a) and magnitude (b)) across seven sub-periods, respectively. Each sub-period is drawn in a different color. Moreover, we purposely highlight the distributions of some sub-periods that behave relatively differently from the rest using solid lines to help readers better locate them in the figures, while other sub-periods are marked by dashed lines. We find that both intra-community and inter-community contagion densities are dramatically amplified during the Covid-19 Outbreak period (solid orange line), compared to other time intervals, as shown in Figures 5(a) and 6(a). This corresponds to one of the most turbulent phases in financial markets, characterized by sharp price declines, heightened volatility, and widespread investor pessimism [82]. This result suggests that the amplified turbulence in financial markets and the extreme reactions of investors (e.g. purchasing and selling herding behavior) caused by the pandemic trigger contagion not only within communities of highly similar assets but also across different asset communities. Furthermore, the Bull Time 2 sub-period (solid red line) also shows a moderate increase in contagion density, although this phenomenon is much weaker than the Covid-19 Outbreak time. This was a period when most assets experienced a price surge, indicating that contagion is not only a feature of negative shocks but can also be triggered under favorable market conditions. In fact, the Covid-19 pandemic had a much stronger impact on financial markets than the bullish time, with widespread systemic disruptions and heightened investor attention [12], explaining the more pronounced contagion effect observed during that time.
@@
Figure 5: Distributions of a) intra-community contagion densities and b) intra-community contagion magnitudes during each of seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia 1 and Ukraine-Russia 2. We use solid lines to highlight the distributions of notable sub-periods. Otherwise, they are drawn by dashed lines
@@
Figure 6: Distributions of a) inter-community contagion densities and b) inter-community contagion magnitudes during each of seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia 1 and Ukraine-Russia 2. We use solid lines to highlight the distributions of notable sub-periods, otherwise, they are drawn by dashed lines
@@
Remarkably, the distributions of contagion magnitudes across all sub-periods remain relatively similar, indicating a consistent level of magnitude over time, regardless of market conditions. This pattern holds true for both intra-community and inter-community contagions, suggesting that while the number of contagion links tends to increase during turbulent times, the average strength of these links remains stable. This phenomenon seems to be attributed to the portfolio diversification and risk-hedging strategy from investors in which they adjust their holdings across multiple assets rather than only focus on a particular one, especially during a financial crisis [110]. These widespread but diluted approaches can raise the scale of contagion without amplifying the strength of individual spillovers.
To reinforce this result, we introduce two contagion metrics based on the measures defined in Subsection 4.6, including Total Contagion Density (TCD) and Total Contagion Magnitude (TCM). These metrics allow us to quantitatively measure and compare contagion density and magnitude across sub-periods. Specifically, for each sub-period, we compute the total contagion by summing the average intra-community and inter-community contagion, separately for density and magnitude. Details are as follows:
  1. 1.
    Total Contagion Density (TCD): to measure the total average contagion density within a given period of time by considering both intra-community and inter-community contagion.
    Let $SD_{Intra},SD_{Inter}$ be sets of intra-community and inter-community contagion densities within a period of time, respectively. Let $n_{SD},m_{SD}$ be the cardinality of $SD_{Intra},SD_{Inter}$, respectively.
    $TCD=\frac{\sum_{i\in SD_{Intra}}i}{n_{SD}}+\frac{\sum_{j\in SD_{Inter}}j}{m_{SD}}$ (16)
  2. 2.
    Total Contagion Magnitude (TCM): to measure the total average contagion magnitude within a given period of time by considering both intra-community and inter-community contagion.
    Let $SM_{Intra},SM_{Inter}$ be sets of intra-community and inter-community contagion magnitudes within a period of time, respectively. Let $n_{SM},m_{SM}$ be the cardinality of $SM_{Intra},SM_{Inter}$, respectively.
    $TCM=\frac{\sum_{i\in SM_{Intra}}i}{n_{SM}}+\frac{\sum_{j\in SM_{Inter}}j}{m_{SM}}$ (17)
Table 5 reports the total contagion density and magnitude for each sub-period. As can be seen, the TCD value during the Covid-19 Outbreak is substantially higher than in any other sub-periods, reaching 59.28, whereas most other sub-periods fall within the range of 22 to 26. Additionally, the results also indicate that the Bull Time 2 sub-period exhibits slightly elevated total contagion density compared with the remaining periods, although still far below the Covid-19 Outbreak level. Regarding the total contagion magnitude, although the Covid-19 Outbreak again records the highest TCM value, its difference from other sub-periods is relatively small since most of them show relatively similar levels. To this end, these findings support the patterns observed in Figures 6 and 7: contagion density tends to amplify during turbulent periods while contagion magnitude appears much less sensitive to market conditions.
Table 5: Total contagion density and magnitude in each sub-period
Measure Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
Total Contagion Magnitude 0.336 0.596 0.399 0.565 0.547 0.472 0.499
Total Contagion Density 22.236 59.278 24.493 30.339 26.577 22.414 22.079
@@
We now compare the contagion effects between intra-community and inter-community during each sub-period. Utilizing the distribution of inter-community contagions described above (shown in Figure 6), our idea for this experiment is to compare each intra-community contagion with the distribution to see if the intra-community one behaves similarly to the inter-community contagions or not. In this regard, an intra-community contagion is considered to be significantly stronger than inter-community contagions within a sub-period if its quantile value falls within the right-hand side rejection area in the inter-community contagion distribution, with the significance level being 1% (i.e. the quantile value is greater than or equal to the 99th). Otherwise, it behaves similarly to inter-community contagions. For this, Tables 6-7 show the contagion density and magnitude, respectively, within each community along with their quantiles (inside the brackets) in the corresponding inter-community contagion distribution for each of the seven sub-periods. Generally, both tables reveal that the intra-community contagion effects do not look much different from inter-community contagion effects since most intra-community contagion signals are similar to inter-community contagion signals. Specifically, the density (Table 6) and magnitude (Table 7) values within each community mostly belong to the 99% confidence interval of their corresponding inter-community contagion distribution and are outside the rejection areas, indicating that intra-community contagion tends to follow the same behavior as inter-community contagion, across different time intervals and market conditions. By contrast, only a few communities in certain sub-periods reveal an abnormal behavior where their intra-community contagion densities or magnitudes are significantly stronger than the others and are off the inter-community contagion distributions. This phenomenon is more frequent for contagion density. For example, community #9 shows a contagion density exceeding 26% in Bull Time 1 and 32% in Ukraine-Russia Conflict 1, placing it at the 100th percentile of the corresponding inter-community contagion density distributions for those periods. Similarly, communities #7 (100th percentile) and #11 (99th percentile) in Bull Time 2 also exhibit unusually high contagion densities. As for contagion magnitude, only communities #1 and #14 in the Ukraine-Russia Conflict 2 sub-period surpass inter-community levels. A common characteristic of these communities is that the assets within each community mostly come from the same business sector, either Information Technology, Financials or cryptocurrency. For this result, a possible reason is that these assets are closely relevant to each other, sharing common fundamental drivers such as macroeconomic variables, regulatory changes and investors’ behaviors. This similarity in risk exposures and underlying characteristics naturally leads to higher correlation in price movements, which tends to trigger contagion within the group, especially during market uncertainties [3]. However, these communities only account for a very small part of the network.
@@
To further assess the similarity between intra-community and inter-community contagion effects, we apply the Kolmogorov–Smirnov test [111], which evaluates whether two samples are drawn from the same underlying distribution. A statistically significant result indicates that the two distributions differ, whereas a p-value greater than the chosen significance level suggests that the samples are statistically indistinguishable and therefore similar in their characteristics. In this present study, we set the confidence level to be 99%, which is equivalent to a significance level of 0.01. P-values of the Kolmogorov-Smirnov tests between intra-community and inter-community contagion effects during different sub-periods are shown in Table 8.
Table 8: P-values of the Kolmogorov–Smirnov tests comparing intra-community and inter-community contagion in each sub-period. A p-value below 1% (0.01) indicates that the two distributions differ, whereas a p-value above this threshold suggests no statistical difference between them. Contagion is evaluated in two aspects: density and magnitude. A star symbol denotes tests with p-values below 0.01.
Contagion
Measure
Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
Density 0.272 0.210 0.513
0.002*
0.023 (adjusted)
0.826 0.402 0.052
Magnitude 0.107 0.475 0.589 0.069 0.070 0.236 0.272
Across all sub-periods, almost all p-values exceed 0.01, indicating no statistically significant difference between intra-community and inter-community contagion in both density and magnitude. The only exception occurs in the Bull Time 2 sub-period, where the contagion density between intra-community and inter-community shows a statistically significant difference, with p-values below 0.01. A closer examination reveals that this result is driven by only a few communities exhibiting relatively high contagion density, as mentioned and explained earlier. Once these communities are removed from the analysis, the p-value rises above the significance threshold, confirming the similarity between intra-community and inter-community contagion. Thus, even during this sub-period, the majority of communities display contagion patterns consistent with those observed across communities.
Another noteworthy finding is that although cryptocurrencies are highly correlated with each other and exhibit distinct behavior from other traditional assets, as shown in Subsection 5.1, the contagion density and magnitude within communities of cryptocurrencies generally do not differ from other asset groups. The only notable exception is during the Bull Time 2 sub-period, when the contagion density among cryptocurrencies stood at the highest level across the entire network, exceeding both intra- and inter-community contagion densities. As discussed earlier, cryptocurrencies share several common fundamental drivers, especially the strong sensitivity of investors to positive market conditions [12], combined with the positive momentum of financial markets during this bullish time, which explains this high contagion density.
Overall, the above results suggest that financial contagion tends to manifest similarly both within and across communities since there is no significant difference between intra-community and inter-community contagions. This implies that contagion is unlikely to be an important factor in the formation of communities. Supporting this point is the case of cryptocurrencies: although they exhibit strong internal correlations and form distinct communities, their contagion levels are generally comparable to those of other asset communities and cross-communities throughout most periods.
@@ 5.3 Directional Inter-community Contagion
In this subsection, we take a closer look at inter-community contagion effects by analyzing how shocks propagate directionally between different communities. Specifically, we investigate how one community transmits shocks to another and how it, in turn, receives shocks. Our goal is to identify whether there exist communities acting as dominant transmitters (e.g. showing strong outgoing contagion to a high number of communities) or receivers (showing strong incoming contagion from a high number of communities). Also, as discussed in the previous subsection, contagion magnitude shows rather small variation over time and offers few meaningful insights. Therefore, for the sake of clarity and simplicity, we focus solely on contagion density in this part of the analysis. To this end, we use the directional contagion density metric described in Equation 14 for this experiment.
@@
Figure 7: Heatmap of directional inter-community contagion density (a) and its corresponding graph representation of top 25% in contagion density (b). In both sub-figures, a darker color indicates a higher value. An arrow represents the direction of contagion from one community to another.
Indeed, our experimental results show that there are always communities acting as strong transmitters in each sub-period, spreading shocks to many other communities with significantly higher contagion density levels relative to the overall network. On the other hand, this pattern for the case of receiving communities is not observed across all sub-periods, suggesting that there is no community intensively receiving shocks from many others. Notably, communities that act as strong transmitters in the network are often composed of stocks and US ETFs dominated by holdings in the Information Technology sector. Besides, sectors like Healthcare, Financials, Energy and cryptocurrencies also emerge as strong transmitters in specific periods. To support these findings, we present two examples: the graph visualizations of the top 25% inter-community contagions in terms of contagion density (Figures 8(a) and 8(b)) and their quantitative details (Tables 9 and 9) for the Bull Time 3 and Ukraine-Russia Conflict 1 sub-periods, respectively. Specifically, Table 9 shows that communities #1 and #2 during the Bull Time 3 sub-period act as dominant transmitters, transmitting high-density shocks (i.e. belong to the top 25% of inter-community contagions in terms of the contagion density) to 14 and 12 out of 16 communities, respectively. These numbers are significantly higher than those for other communities. Moreover, the average contagion densities from communities #1 and #2 to other communities also stand at the highest levels, with 23.75% and 20.10%, respectively. Consequently, these two communities are highlighted in the darkest blue in Figure 8(a). Notably, community #1 is composed predominantly of stocks and US ETFs dominated by holdings in Information Technology (75%) while Healthcare assets contribute approximately 86% to community #2. Likewise, during the Ukraine-Russia Conflict 1 sub-period, communities #6, #9 and #10 emerge as dominant transmitters. For this, Financials assets primarily make up community #9, whereas Information Technology assets account for the majority in the other two communities. Further details are provided in Table 9. Beyond these examples, a summary of the dominant transmitters in each sub-period and their primary business sectors is provided in the supporting file S4.
@@
Figure 8: Graph representation of the top 25% inter-community contagions in terms of the contagion density during the Bull Time 3 (a) and Ukraine-Russia Conflict 1 (b) sub-periods. A darker color indicates a higher value. An arrow represents the direction of contagion from one community to another.
Table 9: Quantitative details of the top 25% inter-community contagion links in terms of density during the (a) Bull Time 3 and (b) Ukraine-Russia Conflict 1 sub-periods. Com represents the community, #In is the number of incoming contagion links to a community, #Out is the number of outgoing contagion links from a community, Avg_In is the average contagion density of incoming contagion links to a community, and Avg_Out is the average contagion density of outgoing contagion links from a community. Communities acting as dominant transmitters are highlighted in bold.
@@
The strong influence of these sectors can be attributed to their intrinsic characteristics and also the prevailing market conditions during the examined periods. In particular, Information Technology assets consistently attract significant investor attention due to their strong growth potential and dominant market capitalization in the financial markets, making them highly influential in shaping both investor behavior and overall market performance. As a result, price movements in this sector often have a broad impact across various business sectors and types of investment vehicle. In fact, the sector’s influence has grown even more remarkably since 2020 thanks to the accelerated global technology adoption followed by the boom in generative AI [112, 113]. Regarding other aforementioned sectors, their strong influence appears to be more context-dependent. For instance, during the pandemic, the Healthcare sector gained prominence as companies were at the forefront of efforts to combat Covid-19, including the production of protective equipment and vaccines [114, 115]. This rapid development elevated their importance and impact on both financial markets and the broader economy. Likewise, the Financials sector plays a vital role in reflecting and shaping economic conditions [116], thereby carrying a strong influence in financial markets, especially during periods of uncertainty like the Ukraine-Russia conflict. To this end, our results are supported to some extent by existing studies, although they are not directly relevant to our present topic and dataset context. For instance, Giudici and Parisi in [117] examined the systemic risk among ten countries using Corporate Default Swap (CDS) spreads over a period from 2006 to 2016. By introducing a novel approach that integrates partial correlations and correlation networks into VAR models, they discovered a clustering effect between the countries in terms of credit risk contagion. Specifically, they found that some countries tend to act as risk importers, while some often behave as risk exporters. Despite the differences in the two topics, this finding shares some common ground with the results of our experiments.
@@
6 Robustness Analyses
In this section, we run robustness analyses of our results when using other approaches in conducting experiments. We cover a wide range of experiments to prove that our results are consistent even when different approaches in different parts of our experiments are selected. These include: 1) evaluating the community structure (i.e. asset partition) using an alternative random matrix theory to filter noise signals; 2) reassessing contagion results under a stricter 99% confidence level; 3) examining the community structure derived from an alternative network construction—the Planar Maximally Filtered Graph (PMFG).
6.1 Marchenko-Pastur Random Matrix Theory for Noise Filtering
Our noise filtering procedure in this study relies on the Tracy-Widom random matrix theory. Since the purpose of this preprocessing step is to find the asset partition (or community structure) without the effect of noise in financial markets, we evaluate the consistency of the resulting partition with that obtained using the Marchenko–Pastur random matrix theory [26, 68]. In this regard, we adopt the v-measure metric [26] to assess the similarity between two community structures obtained by the two aforementioned methods. This metric ranges from 0 to 1, where a value of 1 indicates identical community structures while a value of 0 indicates completely dissimilar ones. Results for this comparison are shown in Table 10. Additionally, we also investigate the number of eigenvalues considered to be informative under each method to determine whether their selections differ. This result is displayed in Table 11.
Table 10: Community structure comparison between Tracy-Widom-based noise filtering and Marchenko-Pastur-based noise filtering.
Measure Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
v-measure 0.90 1.00 0.94 0.82 0.87 0.90 0.82
Table 11: Number of informative eigenvalues in the correlation matrix for each sub-period, as determined by the Tracy–Widom and Marchenko–Pastur methods.
Sub-period Tracy-Widom Marchenko-Pastur
Pre-Covid-19 18 15
Covid-19 Outbreak 8 8
Bull Time 1 15 13
Bull Time 2 17 14
Bull Time 3 16 13
U-R Conflict 1 13 11
U-R Conflict 2 16 13
As shown in Table 10, the community structures obtained between two theories exhibit a high degree of similarity, with v-measure values exceeding 80% in all sub-periods and most above 90%. This indicates that the use of either Tracy-Widom or Marchenko-Pastur random matrix theory does not cause significant changes in the noise filtering result. However, looking at Table 11, we notice that the number of informative eigenvalues identified by Tracy-Widom theory tends to be higher than that obtained using the other. This occurs because Marchenko-Pastur relies on a strict theoretical formula to filter noise in the correlation matrix, making it sensitive to the sample size, as discussed earlier in Subsection 2.3. As a result, some eigenvalues that carry meaningful information may still fall below the Marchenko-Pastur threshold due to their relatively small magnitude and are therefore classified as noise. In contrast, Tracy-Widom is less sensitive to data size and its threshold is based on empirical experimentation rather than a fixed theoretical distribution. This allows the theory to adapt more flexibly to the data and retain additional informative eigenvalues.
It is worth noting that the Marchenko-Pastur random matrix theory has been widely adopted in previous studies because it preserves the most significant information while effectively filtering out noise. However, the more flexible and empirically driven nature of the Tracy-Widom one allows it to capture additional structure in the correlation matrix that Marchenko–Pastur may overlook by discarding small but informative eigenvalues.
6.2 Contagion Coefficients Measured at the 99% Confidence Level
We are currently employing a 90% confidence level to filter out unreliable contagion coefficients estimated from the VAR model, as discussed in Subsection 4.5. Recall that each contagion coefficient in this study measures the extent to which movements in one asset influence another. Therefore, coefficients equal to 0.0 or non-zero coefficients that fail to meet the 90% confidence threshold indicate no contagion relationship between the corresponding pair of assets. To test the robustness of the original contagion coefficients (i.e. at the 90% confidence level), we evaluate them using a stricter confidence level of 99%. Table 12 reports the percentage of original non-zero contagion coefficients that are statistically significant at the 90% level and still remain significant at the 99% level.
Table 12: Percentage of the original non-zero contagion coefficients used in this study that are still statistically significant at the 99% confidence level.
Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
90.07 94.96 91.17 91.29 91.38 90.84 90.97
Across all sub-periods, more than 90% of the contagion coefficients that are statistically significant at the 90% confidence level remain significant when evaluated at the stricter 99% threshold. This implies that, in each sub-period, at most 10% of the original coefficients are filtered out under the higher confidence level. Indeed, since each sub-period is sufficiently long for reliable VAR estimation, the resulting contagion coefficients generally have very low p-values, generally close to 0.0. Consequently, increasing the confidence level has a low impact on our original contagion coefficients.
To assess the robustness of our original findings under a stricter confidence threshold, we repeat the key experiments using the new 99% confidence level and compare the new results with our original ones. These include: 1) examining the contagion effects within and between communities across all sub-periods to assess the evolution of contagion over time; 2) comparing intra-community and inter-community contagion; 3) analyzing directional contagions between communities. Overall, the main findings remain unchanged. However, contagion densities both within and between communities tend to be slightly lower than before. This decline is expected since raising the confidence level reduces the number of reliable contagion coefficients (i.e. the contagion links between assets). By contrast, the new contagion magnitudes do not exhibit this reduction. Details of these experiments are provided in the supporting file S5.
6.3 Planar Maximally Filtered Graph (PMFG) for Network Construction
In our study, we first construct an asset network using the Minimum Spanning Tree (MST) and then apply the Louvain community detection algorithm to identify its community structure. However, because the MST retains only N-1 edges, it may discard many potentially important relationships between financial assets. This reduction can bias the resulting communities and, in turn, lead to unreliable community-based experimental results. To address this concern, we examine the robustness of the MST-derived community structure and its relevant contagion findings by comparing them with those obtained from another well-established network representation: the Planar Maximally Filtered Graph (PMFG) [118]. In brief, this graph preserves significantly more information by maintaining planarity - meaning no edges intersect - while still retaining the strongest asset relationships. As a result, it offers a richer and more informative network with 3(N-2) edges. Figure 9(a) presents the PMFG-based network and its corresponding community structure for the Pre–Covid-19 sub-period as an example. The sector distribution of each identified community is then displayed in Figure 9(b).
(a) Network and Community Structure
(b) Sector Distribution
Figure 9: Network and Community structure (a) and sector distribution per community (b) of 221 assets during the Pre-Covid-19 period, based on the Planar Maximally Filtered Graph (PMFG). A US ETF is assigned to a specific business sector if it is designed to track the performance of stocks within that sector. Otherwise, it is unclassified and labeled as US ETF (i.e. US ETFs that do not track a specific sector or those that include stocks from multiple sectors.
We begin by comparing the similarity between community structures derived from our original MST-based approach and those obtained using the PMFG, using the v-measure as the evaluation metric. Results are presented in Table 13, showing that the PMFG-based communities are largely consistent with those from the MST. Specifically, across all sub-periods and market conditions, the similarity remains high - averaging around 80%, indicating that the community structure is relatively stable regardless of the underlying network construction method. Notably, we observe that the number of communities identified using the PMFG is consistently lower than that from the MST. This phenomenon arises from the denser network structure produced by the PMFG, which includes more connections between assets. As a result, each community tends to contain more assets. Table 14 shows the number of communities in each sub-period, as detected from the MST and PMFG, respectively.
Table 13: Community structure comparison between MST and PMFG
Measure Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
v-measure 0.85 0.76 0.79 0.79 0.78 0.84 0.85
Table 14: Number of communities in each sub-period, as detected from MST and PMFG.
Sub-period MST-based PMFG-based
Pre-Covid-19 15 10
Covid-19 Outbreak 13 8
Bull Time 1 16 9
Bull Time 2 16 10
Bull Time 3 17 9
U-R Conflict 1 15 9
U-R Conflict 2 15 11
To validate our original findings, we rerun the main contagion analyses similarly to the previous Subsection 6.2 using the community structure derived from the PMFG. Results from all experiments remain consistent with those obtained using the MST-based structure, confirming the robustness of our conclusions. To this end, although the PMFG yields a slightly different community structure as reflected in the differing number of communities, the main findings reported in this study are fully preserved. Detailed results of these experiments are provided in the supporting file S6.